# The Jaynes-Cummings model as an autonomous Maxwell demon

### v2

@@ III.1 Measurement by the cavity
$\hat{T}_{eg}=\hat{T}_{ge}^{\dagger}$ $=\frac{1}{2}|e\rangle\langle g|\otimes[-\hat{c}_{1}\hat{\rho}_{C}\left(0\right)\hat{s}_{0}^{\dagger}+\hat{s}_{1}\hat{\rho}_{C}\left(0\right)\hat{c}_{0}-i\hat{c}_{1}\hat{\rho}_{C}\left(0\right)\hat{c}_{0}-i\hat{s}_{1}\hat{\rho}_{C}\left(0\right)\hat{s}_{0}^{\dagger}]$ (S38)
$=\frac{-i}{2}|e\rangle\langle g|\otimes\left(\hat{c}_{1}+i\hat{s}_{1}\right)\hat{\rho}_{C}\left(0\right)\left(\hat{c}_{0}-i\hat{s}_{0}^{\dagger}\right)$
$=\frac{-i}{2}|e\rangle\langle g|\otimes\left(\hat{c}_{1}+i\sqrt{n_{0}}\hat{s}_{1}^{{}^{\prime}}\right)\hat{\rho}_{C}\left(0\right)\left(\hat{c}_{0}-i\sqrt{n_{0}}\hat{s}_{0}^{{}^{\prime}}\right)$
$=-\frac{ie^{\frac{igt}{2\sqrt{n_{0}}}}}{2}|e\rangle\langle g|\otimes e^{\frac{igt\hat{n}}{2\sqrt{n_{0}}}}\hat{\rho}_{C}\left(0\right)e^{-\frac{igt\hat{n}}{2\sqrt{n_{0}}}}+O\left(\frac{1}{n_{0}}\right)$
@@
$\hat{c}_{0}+i\sqrt{n_{0}}\hat{s}_{0}^{{}^{\prime}}$ $=\cos\left(gt\sqrt{\hat{n}}\right)+i\sqrt{n_{0}}\frac{\sin\left(gt\sqrt{\hat{n}}\right)}{\sqrt{\hat{n}}}$ (S39)
@@
$\hat{T}_{gg}$ $=\frac{1}{2}|g\rangle\langle g|\otimes e^{\frac{igt\hat{n}}{2\sqrt{n_{0}}}}\hat{\rho}_{C}\left(0\right)e^{-\frac{igt\hat{n}}{2\sqrt{n_{0}}}}+{\cal O}\left(\frac{1}{n_{0}}\right)$ (S40)
@@
$\hat{U}\left(t\right)|+_{y}\rangle\langle+_{y}|\otimes\hat{\rho}_{C}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ $=\frac{1}{2}\left[\begin{matrix}1&-ie^{\frac{igt}{2\sqrt{n_{0}}}}\\ ie^{-\frac{igt}{2\sqrt{n_{0}}}}&1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)+{\cal O}\left(\frac{1}{n_{0}}\right)$ (S41)
@@
$\hat{U}\left(t\right)|-_{y}\rangle\langle-_{y}|\otimes\hat{\rho}_{C}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ $=\frac{1}{2}\left[\begin{matrix}1&ie^{-\frac{igt}{2\sqrt{n_{0}}}}\\ -ie^{\frac{igt}{2\sqrt{n_{0}}}}&1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)$ (S42)
$\hat{U}\left(t\right)|+_{y}\rangle\langle-_{y}|\otimes\hat{\rho}_{C}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ $=\frac{e^{igt\sqrt{n_{0}}}}{2}\left[\begin{matrix}e^{\frac{igt}{\sqrt{n_{0}}}}&ie^{\frac{igt}{2\sqrt{n_{0}}}}\\ ie^{\frac{igt}{2\sqrt{n_{0}}}}&-1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)$ (S43)
$\hat{U}\left(t\right)|-_{y}\rangle\langle+_{y}|\otimes\hat{\rho}_{C}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ $=\frac{e^{-igt\sqrt{n_{0}}}}{2}\left[\begin{matrix}e^{-\frac{igt}{\sqrt{n_{0}}}}&-ie^{-\frac{igt}{2\sqrt{n_{0}}}}\\ -ie^{\frac{-igt}{2\sqrt{n_{0}}}}&-1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)$ (S44)
@@
$\hat{\rho}_{QC}\left(t\right)$ $=$ $\hat{U}\left(t\right)\sum_{\nu,\mu=\pm}\rho_{\nu,\mu}|\nu_{y}\rangle\langle\mu_{y}|\otimes D\left(\sqrt{n_{0}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\sqrt{n_{0}}\right)\hat{U}^{\dagger}\left(t\right)$ (S47)
$=$ $\frac{\rho_{++}}{2}\left[\begin{matrix}1&-ie^{\frac{igt}{2\sqrt{n_{0}}}}\\ ie^{-\frac{igt}{2\sqrt{n_{0}}}}&1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)$
$+\frac{\rho_{--}}{2}\left[\begin{matrix}1&ie^{\frac{-igt}{2\sqrt{n_{0}}}}\\ -ie^{\frac{igt}{2\sqrt{n_{0}}}}&&1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)+$
$+\frac{\rho_{+-}e^{igt\sqrt{n_{0}}}}{2}\left[\begin{matrix}e^{\frac{igt}{\sqrt{n_{0}}}}&ie^{\frac{igt}{2\sqrt{n_{0}}}}\\ ie^{\frac{igt}{2\sqrt{n_{0}}}}&&-1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)$
$+\frac{\rho_{-+}e^{-igt\sqrt{n_{0}}}}{2}\left[\begin{matrix}e^{-\frac{igt}{\sqrt{n_{0}}}}&-ie^{-\frac{igt}{2\sqrt{n_{0}}}}\\ -ie^{\frac{-igt}{2\sqrt{n_{0}}}}&-1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)+O\left(\frac{1}{n_{0}}\right)$
$=$ $\rho_{++}|+_{y}\left(t\right)\rangle\langle+_{y}\left(t\right)|\otimes\hat{\rho}_{C}^{++}\left(t\right)+\rho_{--}|-_{y}\left(t\right)\rangle\langle-_{y}\left(t\right)|\otimes\hat{\rho}_{C}^{--}\left(t\right)$
$+\rho_{+-}e^{igt\sqrt{n_{0}}+\frac{igt}{\sqrt{n_{0}}}}|+_{y}\left(t\right)\rangle\langle-_{y}\left(t\right)|\otimes\hat{\rho}_{C}^{+-}\left(t\right)++\rho_{-+}e^{-igt\sqrt{n_{0}}-\frac{igt}{\sqrt{n_{0}}}}|-_{y}\left(t\right)\rangle\langle+_{y}\left(t\right)|\otimes\hat{\rho}_{C}^{-+}\left(t\right)$
@@
where |±$|\pm_{y}\rangle\left(t\right)=\frac{1}{\sqrt{2}}\left(\begin{matrix}1\\ \pm ie^{\mp\frac{igt}{2\sqrt{n_{0}}}}\end{matrix}\right)$, and $\hat{\rho}_{C}^{\nu\mu}\left(t\right)$ is defined in Eq. (2) of main text.
We now show that the trace of the terms $\hat{\rho}_{C}^{\pm\mp}\left(t_{c}\right)$ becomes negligible at $t_{c}$. We have:
$\text{Tr}[\hat{\rho}_{C}^{\pm\mp}\left(t\right)]$ $=e^{\pm igt(\sqrt{n_{0}}+1/\sqrt{n_{0}})}\text{Tr}[D\left(\sqrt{n_{0}}e^{\pm i\frac{gt}{2\sqrt{n_{0}}}}\right)w_{\beta\left(0\right)}D^{\dagger}\left(\sqrt{n_{0}}e^{\mp i\frac{gt}{2\sqrt{n_{0}}}}\right)]$ (S48)
$=e^{\pm igt(\sqrt{n_{0}}+1/\sqrt{n_{0}})}e^{\pm in_{0}\sin\frac{gt}{\sqrt{n_{0}}}}\text{Tr}[w_{\beta\left(0\right)}D\left(\pm 2i\sqrt{n_{0}}\sin\left(\frac{gt}{2\sqrt{n_{0}}}\right)\right)]$
$=e^{\pm igt(\sqrt{n_{0}}+1/\sqrt{n_{0}})}\left(1-e^{-\beta\left(0\right)}\right)e^{\pm in_{0}\sin{\frac{gt}{\sqrt{n_{0}}}}}\frac{1}{\pi}\int d^{\left(2\right)}\alpha\langle\alpha|e^{-\beta\left(0\right)a^{\dagger}a}D\left(\pm 2i\sqrt{n_{0}}\sin\left(\frac{gt}{2\sqrt{n_{0}}}\right)\right)|\alpha\rangle$
@@
$\text{Tr}[\hat{\rho}_{C}^{\pm\mp}\left(t\right)]$ $=e^{\pm igt(\sqrt{n_{0}}+1/\sqrt{n_{0}})}e^{\pm in_{0}\sin{\frac{gt}{\sqrt{n_{0}}}}}e^{-2n_{0}\sin^{2}{\frac{gt}{2\sqrt{n_{0}}}}}e^{-\frac{4n_{0}\bar{n}}{\left(1+\bar{n}\right)^{2}}\sin^{2}{\frac{gt}{2\sqrt{n_{0}}}}}$ (S49)
@@
$\text{Tr}[\hat{\rho}_{C}^{\pm\mp}\left(t\right)]$ $=e^{\pm 2igt\sqrt{n_{0}}}e^{-\frac{g^{2}t^{2}}{2}}e^{-\frac{g^{2}t^{2}\bar{n}}{\left(1+\bar{n}\right)^{2}}+\frac{\bar{n}}{\left(1+\bar{n}\right)^{2}}\frac{g^{4}t^{4}}{12n_{0}}},$ (S50)
which tends to 00.
@@
$\hat{\rho}_{Q}\left(t_{c}\right)$ $=\rho_{++}|+_{y}\left(t_{c}\right)\rangle\langle+_{y}\left(t_{c}\right)|+\rho_{--}|-_{y}\left(t_{c}\right)\rangle\langle-_{y}\left(t_{c}\right)|+O\left(\frac{1}{n_{0}}\right)$ (S51)
@@
$\hat{\rho}_{Q}\left(t_{c}\right)$ $=\rho_{++}|+_{y}\rangle\langle+_{y}|+\rho_{--}|-_{y}\rangle\langle-_{y}|+O\left(\frac{1}{n_{0}}\right)$ (S52)
III.2 Autonomous feedback

### v4

@@ C.1 Perturbative expansion of the global atom–cavity state
$\hat{T}_{eg}=\hat{T}_{ge}^{\dagger}$ $=\frac{1}{2}|e\rangle\langle g|\otimes[-\hat{c}_{1}\hat{\rho}_{\text{C}}\left(0\right)\hat{s}_{0}^{\dagger}+\hat{s}_{1}\hat{\rho}_{\text{C}}\left(0\right)\hat{c}_{0}-i\hat{c}_{1}\hat{\rho}_{\text{C}}\left(0\right)\hat{c}_{0}-i\hat{s}_{1}\hat{\rho}_{\text{C}}\left(0\right)\hat{s}_{0}^{\dagger}]$ (68)
$=\frac{-i}{2}|e\rangle\langle g|\otimes\left(\hat{c}_{1}+i\hat{s}_{1}\right)\hat{\rho}_{\text{C}}\left(0\right)\left(\hat{c}_{0}-i\hat{s}_{0}^{\dagger}\right)$
$=\frac{-i}{2}|e\rangle\langle g|\otimes\left(\hat{c}_{1}+i\sqrt{n_{0}}\hat{s}_{1}^{{}^{\prime}}\right)\hat{\rho}_{\text{C}}\left(0\right)\left(\hat{c}_{0}-i\sqrt{n_{0}}\hat{s}_{0}^{{}^{\prime}}\right)$
$=-\frac{ie^{\frac{igt}{2\sqrt{n_{0}}}}}{2}|e\rangle\langle g|\otimes e^{\frac{igt\hat{n}}{2\sqrt{n_{0}}}}\hat{\rho}_{\text{C}}\left(0\right)e^{-\frac{igt\hat{n}}{2\sqrt{n_{0}}}}+O\left(\frac{1}{n_{0}}\right)$
@@
$\hat{c}_{0}+i\sqrt{n_{0}}\hat{s}_{0}^{{}^{\prime}}$ $=\cos\left(gt\sqrt{\hat{n}}\right)+i\sqrt{n_{0}}\frac{\sin\left(gt\sqrt{\hat{n}}\right)}{\sqrt{\hat{n}}}$ (69)
@@
$\hat{T}_{gg}$ $=\frac{1}{2}|g\rangle\langle g|\otimes e^{\frac{igt\hat{n}}{2\sqrt{n_{0}}}}\hat{\rho}_{\text{C}}\left(0\right)e^{-\frac{igt\hat{n}}{2\sqrt{n_{0}}}}+{\cal O}\left(\frac{1}{n_{0}}\right)$ (70)
@@
$\hat{U}\left(t\right)|+_{y}\rangle\langle+_{y}|\otimes\hat{\rho}_{\text{C}}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ $=\frac{1}{2}\left[\begin{matrix}1&-ie^{\frac{igt}{2\sqrt{n_{0}}}}\\ ie^{-\frac{igt}{2\sqrt{n_{0}}}}&1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)+{\cal O}\left(\frac{1}{n_{0}}\right)$ (71)
@@
$\hat{U}\left(t\right)|-_{y}\rangle\langle-_{y}|\otimes\hat{\rho}_{\text{C}}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ $=\frac{1}{2}\left[\begin{matrix}1&ie^{-\frac{igt}{2\sqrt{n_{0}}}}\\ -ie^{\frac{igt}{2\sqrt{n_{0}}}}&1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)$ (72)
$\hat{U}\left(t\right)|+_{y}\rangle\langle-_{y}|\otimes\hat{\rho}_{\text{C}}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ $=\frac{e^{igt\sqrt{n_{0}}}}{2}\left[\begin{matrix}e^{\frac{igt}{\sqrt{n_{0}}}}&ie^{\frac{igt}{2\sqrt{n_{0}}}}\\ ie^{\frac{igt}{2\sqrt{n_{0}}}}&-1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)$ (73)
$\hat{U}\left(t\right)|-_{y}\rangle\langle+_{y}|\otimes\hat{\rho}_{\text{C}}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ $=\frac{e^{-igt\sqrt{n_{0}}}}{2}\left[\begin{matrix}e^{-\frac{igt}{\sqrt{n_{0}}}}&-ie^{-\frac{igt}{2\sqrt{n_{0}}}}\\ -ie^{\frac{-igt}{2\sqrt{n_{0}}}}&-1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)$ (74)
@@
$\hat{\rho}_{\text{QC}}\left(t\right)$ $=$ $\hat{U}\left(t\right)\sum_{\nu,\mu=\pm}\rho_{\nu,\mu}|\nu_{y}\rangle\langle\mu_{y}|\otimes D\left(\sqrt{n_{0}}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\sqrt{n_{0}}\right)\hat{U}^{\dagger}\left(t\right)$ (75)
$=$ $\frac{\rho_{++}}{2}\left[\begin{matrix}1&-ie^{\frac{igt}{2\sqrt{n_{0}}}}\\ ie^{-\frac{igt}{2\sqrt{n_{0}}}}&1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)$
$+\frac{\rho_{--}}{2}\left[\begin{matrix}1&ie^{\frac{-igt}{2\sqrt{n_{0}}}}\\ -ie^{\frac{igt}{2\sqrt{n_{0}}}}&&1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)+$
$+\frac{\rho_{+-}e^{igt\sqrt{n_{0}}}}{2}\left[\begin{matrix}e^{\frac{igt}{\sqrt{n_{0}}}}&ie^{\frac{igt}{2\sqrt{n_{0}}}}\\ ie^{\frac{igt}{2\sqrt{n_{0}}}}&&-1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)$
$+\frac{\rho_{-+}e^{-igt\sqrt{n_{0}}}}{2}\left[\begin{matrix}e^{-\frac{igt}{\sqrt{n_{0}}}}&-ie^{-\frac{igt}{2\sqrt{n_{0}}}}\\ -ie^{\frac{-igt}{2\sqrt{n_{0}}}}&-1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)+O\left(\frac{1}{n_{0}}\right)$
$=$ $\rho_{++}|+_{y}\left(t\right)\rangle\langle+_{y}\left(t\right)|\otimes\hat{\rho}_{\text{C}}^{++}\left(t\right)+\rho_{--}|-_{y}\left(t\right)\rangle\langle-_{y}\left(t\right)|\otimes\hat{\rho}_{\text{C}}^{--}\left(t\right)$
$+\rho_{+-}e^{igt\sqrt{n_{0}}+\frac{igt}{\sqrt{n_{0}}}}|+_{y}\left(t\right)\rangle\langle-_{y}\left(t\right)|\otimes\hat{\rho}_{\text{C}}^{+-}\left(t\right)++\rho_{-+}e^{-igt\sqrt{n_{0}}-\frac{igt}{\sqrt{n_{0}}}}|-_{y}\left(t\right)\rangle\langle+_{y}\left(t\right)|\otimes\hat{\rho}_{\text{C}}^{-+}\left(t\right)$
@@
where we have introduced the states:
$|\pm_{y}\left(t\right)\rangle=\frac{1}{\sqrt{2}}\left(\begin{matrix}1\\ \pm ie^{\mp\frac{igt}{2\sqrt{n_{0}}}}\end{matrix}\right),$ (76)
and $\hat{\rho}_{\text{C}}^{\nu\mu}\left(t\right)$ is defined in Eq. (19) of main text.
We now show that the trace of the terms $\hat{\rho}_{\text{C}}^{\pm\mp}\left(t_{c}\right)$ becomes negligible at $t_{c}$. We have:
$\text{Tr}[\hat{\rho}_{\text{C}}^{\pm\mp}\left(t\right)]$ $=e^{\pm igt(\sqrt{n_{0}}+1/\sqrt{n_{0}})}\text{Tr}[D\left(\sqrt{n_{0}}e^{\pm i\frac{gt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\sqrt{n_{0}}e^{\mp i\frac{gt}{2\sqrt{n_{0}}}}\right)]$ (77)
$=e^{\pm igt(\sqrt{n_{0}}+1/\sqrt{n_{0}})}e^{\pm in_{0}\sin\frac{gt}{\sqrt{n_{0}}}}\text{Tr}[\hat{w}_{\beta_{C}\left(0\right)}D\left(\pm 2i\sqrt{n_{0}}\sin\left(\frac{gt}{2\sqrt{n_{0}}}\right)\right)]$
$=e^{\pm igt(\sqrt{n_{0}}+1/\sqrt{n_{0}})}\left(1-e^{-\beta_{C}\left(0\right)}\right)e^{\pm in_{0}\sin{\frac{gt}{\sqrt{n_{0}}}}}\frac{1}{\pi}\int d^{\left(2\right)}\alpha\langle\alpha|e^{-\beta_{C}\left(0\right)a^{\dagger}a}D\left(\pm 2i\sqrt{n_{0}}\sin\left(\frac{gt}{2\sqrt{n_{0}}}\right)\right)|\alpha\rangle$
@@
$\text{Tr}[\hat{\rho}_{\text{C}}^{\pm\mp}\left(t\right)]$ $=e^{\pm igt(\sqrt{n_{0}}+1/\sqrt{n_{0}})}e^{\pm in_{0}\sin{\frac{gt}{\sqrt{n_{0}}}}}e^{-2n_{0}\sin^{2}{\frac{gt}{2\sqrt{n_{0}}}}}e^{-\frac{4n_{0}\bar{n}}{\left(1+\bar{n}\right)^{2}}\sin^{2}{\frac{gt}{2\sqrt{n_{0}}}}}$ (78)
@@
$\text{Tr}[\hat{\rho}_{\text{C}}^{\pm\mp}\left(t\right)]$ $=e^{\pm 2igt\sqrt{n_{0}}}e^{-\frac{g^{2}t^{2}}{2}}e^{-\frac{g^{2}t^{2}\bar{n}}{\left(1+\bar{n}\right)^{2}}+\frac{\bar{n}}{\left(1+\bar{n}\right)^{2}}\frac{g^{4}t^{4}}{12n_{0}}},$ (79)
which tends to 0.
@@
$\hat{\rho}_{\text{Q}}\left(t_{c}\right)$ $=\rho_{++}|+_{y}\left(t_{c}\right)\rangle\langle+_{y}\left(t_{c}\right)|+\rho_{--}|-_{y}\left(t_{c}\right)\rangle\langle-_{y}\left(t_{c}\right)|+O\left(\frac{1}{n_{0}}\right)$ (80)
@@
$\hat{\rho}_{\text{Q}}\left(t_{c}\right)$ $=\rho_{++}|+_{y}\rangle\langle+_{y}|+\rho_{--}|-_{y}\rangle\langle-_{y}|+O\left(\frac{1}{n_{0}}\right)$ (81)
C.2 Autonomous feedback